OpenStudy (anonymous):

What number is midway between 2/5 and 1? 4/5 5/6 7/10

OpenStudy (anonymous):

Find the average: sum the two numbers and divide by two.

OpenStudy (experimentx):

the last one

OpenStudy (hoblos):

1+2/5 = 7/5 (7/5) /2 = 7/10